OpenStudy (anonymous):

Determine if the set Z is a group under the binary operation *: x * y =xy +y. If yes, is it abelian? If not, which conditions fail to hold?

OpenStudy (anonymous):

Check for associativity. Does the following hold \(\forall x,y,z\in\mathbb{Z}\)? \[\begin{align*}(x*y)*z&=x*(y*z)\\ (xy+y)*z&=x*(yz+z)\\ (xy+y)z+z&=x(yz+z)+(yz+z)\\ xyz+yz+z&=xyz+xz+yz+z\end{align*}\]

OpenStudy (anonymous):

Do you know what abelian is? I am not sure but this problem has got me confused.

OpenStudy (anonymous):

Abelian means commutative. does x*y=y*x for all x and y?

OpenStudy (anonymous):

Abelian groups are commutative under our binary operation. Here we have the operator \(\otimes:\mathbb{Z}\to\mathbb{Z}\) s.t. \(x\otimes y=xy+y\). I'll let you determine whether it's a group or not (check our axioms -- closure, identity, inverse, associativity). For some \(x,\, y\in\mathbb{Z}\) we have \(x\otimes y=xy+y\). Now, we also know \(y\otimes x=yx+x=xy+x\). Notice that \(x\otimes y=y\otimes x\) only when \(xy+y=xy+x\Leftrightarrow y=x\) and therefore this does not hold in general.

OpenStudy (anonymous):

-some +all i.e. \(\forall x,y\in\mathbb{Z}\)

OpenStudy (anonymous):

So this is not a group because of associativity?